OpenStudy (4everaddicted2anime):

Medal and Fan. A cone with a radius of 3" has a total area of 24pi sq. in. Find the volume of the cone. A: 24pi B: 15pi C: 12pi

OpenStudy (anitas.):

have you tried doing it? if so what did you get? @4everaddicted2anime

OpenStudy (4everaddicted2anime):

The formula for the total area is L,A. + pi*r^2 L.A. = pi*r*l Total area: 24=9pi + 3pi*l 24=12pi*l I don't know where to go from here?

OpenStudy (anitas.):

yes so then your answer will be 24pi..

OpenStudy (4everaddicted2anime):

I'm confused. I need to find the volume. 24 is the total area

OpenStudy (anitas.):

you already got the volume.. ur confusing ur self

OpenStudy (4everaddicted2anime):

Seriously? Thank you so much

OpenStudy (anitas.):

yes! lol